Agent Orange, a defoliant used on a large scale during the Vietnam war, contains about 2 ppm (m/v) of dioxin, a very toxic substance, as an impurity. How many milliliters of Agent Orange would have to be used in order to place 0.01 g of dioxin in the environment?

1
Expert's answer
2022-02-04T05:26:04-0500

Solution:

1 Parts Per Million = 1.0×10-6 Grams Per Milliliter

Therefore,

2 ppm = 2.0×10-6 g/mL


(0.01 g) × (1 mL / 2.0×10-6 g) = 5000 mL


Answer: 5000 milliliters of Agent Orange would have to be used.